# HydroGEM: A Self Supervised Zero Shot Hybrid TCN Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v1

@@ Abstract
Real-time streamflow monitoring networks generate millions of observations annually, yet maintaining data quality across thousands of remote sensors remains labor-intensive and difficult to scale. We introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a foundation model for continental-scale streamflow quality control. HydroGEM uses a two-stage training approach: self-supervised pretraining on 6.03 million clean sequences from 3,724 USGS stations learns general hydrological representations, followed by fine-tuning with synthetic anomalies injected for detection and reconstruction. A hybrid Temporal Convolutional Network(TCN)-Transformer architecture (containing 14.2M parameters) captures both local temporal patterns and long-range dependencies, while hierarchical normalization handles learning across six orders of magnitude in discharge. On a held-out synthetic test set comprising 799 stations where 18 expert-validated anomaly types are observed, HydroGEM achieves F1 = 0.792 for detection accuracy and 68.7% reconstruction-error reduction, outperforming the strongest baseline by 36.3%. Zero-shot transfer to 100 Environment and Climate Change Canada (ECCC) stations—unseen during training—yields F1 = 0.586, exceeding all baselines and demonstrating cross-national generalization. The model maintains consistent detection across correction magnitudes (1–100%) and aligns with operational seasonal patterns, with peak flagging-rates during winter ice-affected periods matching hydrologists’ correction behavior. Architectural separation between simplified training anomalies and complex physical-space test anomalies confirms that performance reflects learned hydrometric principles rather than pattern memorization. HydroGEM is designed for human-in-the-loop workflows—outputs are quality control suggestions requiring expert review, not autonomous corrections. By automating routine screening while preserving human oversight, HydroGEM offers a scalable approach to expanding hydrologists’ effective reach across growing monitoring networks.
@@ 1 Introduction
Real-time hydrological monitoring networks provide essential observations for water resources management, flood forecasting, ecosystem protection, and climate change adaptation [kratzert2019ungauged, kratzert2022caravan, sterle2024camelschem]. The United States Geological Survey operates more than 10,000 stream gauging stations that collectively produce millions of paired discharge and stage measurements each month, forming the backbone of decision-making across federal, state, and local agencies [kratzert2019towards, sterle2024camelschem]. The value of these observations depends directly on data quality [nearing2022lstm].
However, maintaining quality assurance across thousands of remote sensors in harsh environments is difficult. Sensor drift, equipment faults, ice effects, rating curve changes, clock errors, and transmission problems frequently degrade data [sun2019review, dogo2019survey]. Current operational quality control relies primarily on expert hydrologists who visually inspect time series, reason about discharge-stage relationships, and apply site-specific judgment [nearing2022lstm, goldstein2015peeking]. While this expert-centered workflow delivers high-quality data, it does not scale with expanding networks and rising temporal resolution. The bottleneck is not lack of expertise—it is the mismatch between growing data volumes and limited expert capacity.
@@
Machine learning approaches offer potential solutions but face two fundamental obstacles at continental scale. First, hydrological systems are extremely heterogeneous: discharge spans six orders of magnitude across sites, and flow generation mechanisms vary with climate, physiography, and human modification [kratzert2019towards, kratzert2022caravan, lees2021hydrological]. Models trained on individual sites or small regional datasets fail to generalize to unseen locations. Second, labeled anomaly datasets are scarce. Agency quality flags are not designed as ground truth labels that cleanly separate sensor faults from physical phenomena, and comprehensive labeled datasets across thousands of sites do not exist [dogo2019survey, belay2023unsupervised].
Foundation models trained with self-supervised objectives on massive unlabeled datasets provide a path forward [devlin2019bert, dooley2023forecastpfn]. In weather and climate science, transformers pretrained on reanalysis data now match or exceed numerical models for medium-range forecasting [nguyen2023climax, pathak2022fourcastnet, lam2023graphcast]. In hydrology, multi-basin training has enabled transfer learning to ungauged catchments [kratzert2019towards, kratzert2022caravan]. Yet foundation model principles have not been systematically applied to quality control for in situ sensor networks, where challenges include extreme scale heterogeneity, irregular multi-site time series, real-time processing requirements, and mandatory human oversight.
HydroGEM.
@@
We introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a self-supervised foundation model for continental-scale streamflow quality control. HydroGEM qualifies as a foundation model because it: (1) trains on massive diverse data (3724 sites spanning six orders of magnitude in discharge), (2) uses self-supervised pretraining to learn general hydrological representations, (3) shows zero-shot transfer to unseen sites and countries, and (4) uses a modular architecture that allows task-specific adaptation.
@@
HydroGEM uses a two-stage training approach. Stage 1 pretrains a hybrid TCN-Transformer backbone on 6.03 million clean sequences from 3,724 USGS sites using masked reconstruction. A hierarchical normalization scheme—combining log transforms, site-specific standardization, and explicit scale embeddings—allows learning across extreme magnitude ranges while preserving physically meaningful scale structure [goodfellow2016deep]. Stage 2 fine-tunes the pretrained backbone with a detection head using on-the-fly synthetic anomaly injection rather than curated labels [dooley2023forecastpfn, kim2024enhancing]. The training injector applies simplified corruptions (drift, spikes, flatlines, dropouts, clock shifts) in normalized space to encourage learning of fundamental hydrometric consistency principles rather than memorization of specific anomaly signatures. Clean data preservation exceeds 97%, ensuring the model does not corrupt valid observations.
@@
Evaluation emphasizes generalization across three dimensions. We construct a synthetic test set of 799 USGS sites (completely unseen during training) with expert-validated anomalies that are deliberately more complex than training corruptions, creating a four-axis separation (geographic, mathematical, temporal, parameter) that prevents pattern memorization. On this held-out set, HydroGEM achieves 0.70 F1 for anomaly detection and reduces reconstruction error by 70% relative to injected corruptions—though suggested corrections require expert review before operational use. For real-world validation, we evaluate zero-shot transfer to 100 Environment and Climate Change Canada sites using weak labels derived from operational corrections. Reconstructions show promising alignment with agency corrections in several cases, though direct comparison is limited by differences in correction methodology between USGS and ECCC workflows. This cross-national evaluation tests whether learned representations transfer beyond the training data source [kratzert2019towards, deznabi2024zero].
@@
HydroGEM is designed for human-in-the-loop operational workflows [nearing2022lstm, goldstein2015peeking]. A three-tier flagging system passes high-confidence clean data with minimal review, prioritizes uncertain observations for expert inspection, and provides suggested reconstructions with uncertainty estimates for detected anomalies. Reconstructions are proposed corrections—not automated replacements—and require hydrologist approval before integration into official records. All outputs include provenance information suitable for audit and feedback. The goal is to expand hydrologists’ effective reach by automating routine screening and providing focused decision support, not to replace human expertise.
Contributions
@@
This work advances hydrological quality control through:
@@
    A continental-scale foundation model trained on 3724 USGS sites with 6.67 million sequences—an order of magnitude larger than prior multi-site hydrological studies.
@@
Terminology Clean data: USGS approved records with quality code ‘A’ (Stage 1 training) Masked data: Clean data with random timesteps hidden for self-supervised reconstruction (Stage 1) Corrupted data: Clean data with synthetic anomalies injected (Stage 2 fine-tuning) Anomalous data: Real operational data requiring correction (deployment target) Suggested reconstruction: Model-proposed correction requiring hydrologist approval
2 Related Work
Our work bridges operational hydrological quality control, machine learning for time series anomaly detection, and foundation models in Earth system science. We review each area to position HydroGEM’s contributions.
2.1 Operational Hydrological Quality Control
Traditional quality control in hydrological monitoring combines manual expert review with rule-based automated checks [aquarius2024, schmidt2023saqc]. Commercial systems such as AQUARIUS (Aquatic Informatics) implement configurable range checks, rate-of-change thresholds, and statistical comparison filters, providing standardized workflows and audit trails [aquarius2024]. The open-source SaQC framework advances reproducibility through sequential test configurations and traceable quality flags [schmidt2023saqc].
However, rule-based approaches face fundamental limitations. Fixed thresholds require site-specific calibration and capture only obvious outliers. Context-dependent anomalies—such as ice effects that preserve plausible values but bias discharge [mcmillan2018uncertainty], gradual sensor drift within historical ranges [taormina2015neural], or rating curve shifts after channel-altering floods [addor2017large]—often evade detection. Multivariate relationships between discharge and stage, critical for identifying coupled sensor failures or unit conversion errors, are difficult to encode in rule-based systems. As monitoring networks expand and temporal resolution increases, manual calibration becomes prohibitively expensive.
Our approach addresses these limitations by learning multivariate patterns from continental-scale data, allowing context-aware detection without manual threshold tuning while preserving the human-in-the-loop workflows that agencies require.
2.2 Machine Learning for Time Series Anomaly Detection
Classical anomaly detection methods have been applied to environmental monitoring with mixed success. Statistical approaches (z-score tests, ARIMA residuals), distance-based methods (k-nearest neighbors, isolation forests), and one-class classifiers (one-class SVM, autoencoders) provide baseline capabilities [belay2023unsupervised, shi2023groundwater]. Recent environmental applications include wavelet-based multi-resolution analysis for river stage [nourani2014applications] and comparative evaluations of isolation forests versus one-class SVM for groundwater monitoring, where OCSVM achieved 88% precision on synthetic data [shi2023groundwater].
Deep learning approaches show stronger performance by capturing complex temporal dependencies. LSTM networks and Transformer attention mechanisms excel at modeling long-range patterns in time series [hochreiter1997long, he2019temporal]. Temporal convolutional networks enable efficient parallel processing with exponentially large receptive fields [han2023twawdlstm]. Autoencoders—including variational and LSTM-augmented variants—detect anomalies via reconstruction errors [osman2024variational, lee2024lstm]. Generative adversarial networks improve robustness through learned representations of normal data distributions [chen2025gasn]. Synthetic anomaly generation has emerged as a key technique for training when domain-specific labeled data is unavailable [kim2024enhancing].
Despite these advances, prior work suffers from two critical limitations for continental-scale deployment. First, models typically train and evaluate on individual sites or small regional datasets (often <50<50 sites), failing to capture the extreme hydrological heterogeneity spanning snowmelt-dominated montane systems to ephemeral desert washes to regulated lowland rivers. Transfer learning across such diverse regimes remains largely unexplored in hydrological anomaly detection. Second, most approaches require labeled training data—either manual annotations of anomalous periods or curated datasets separating sensor failures from physical phenomena. Such labels are scarce in operational hydrology, where agency quality flags reflect editorial decisions rather than ground truth [dogo2019survey, belay2023unsupervised].
Our two-stage approach resolves both limitations: self-supervised pretraining on massive unlabeled clean data learns general hydrological representations, while synthetic anomaly injection enables supervised fine-tuning without expensive manual labeling [dooley2023forecastpfn, kim2024enhancing].
2.3 Foundation Models in Earth System Science
Foundation models—large networks pretrained on massive datasets with self-supervised objectives, then transferred to diverse downstream tasks—have transformed Earth system science [nguyen2023climax, pathak2022fourcastnet, lam2023graphcast, bi2023accurate]. ClimaX demonstrated that transformers pretrained on heterogeneous climate reanalysis data with masked token prediction effectively transfer to forecasting, downscaling, and multi-variable prediction across spatiotemporal scales [nguyen2023climax]. Weather foundation models (FourCastNet [pathak2022fourcastnet], GraphCast [lam2023graphcast], Pangu-Weather [bi2023accurate]) now achieve medium-range forecast skill competitive with operational numerical weather prediction at orders of magnitude lower computational cost. Industrial-scale efforts such as IBM/NASA’s Prithvi WxC (2.3B parameters) [prithvi2024wxc] and Oak Ridge’s ORBIT (>100>100B parameters) [wang2024orbit] are extending this paradigm to multi-variable Earth system prediction.
In hydrology, Kratzert and colleagues pioneered multi-basin learning by training LSTM networks across hundreds of catchments using the CAMELS dataset [kratzert2019towards], showing that models learn generalizable runoff generation processes that transfer to ungauged basins. Subsequent work confirmed that diverse training across basins improves predictions despite regional hydrological variation [gauch2021proper]. Remote sensing foundation models pretrained on satellite imagery have shown strong transfer to environmental monitoring tasks [ali2025environmental]. Recent perspectives advocate for foundation models as assistive tools in hydrometeorology when developed with appropriate domain constraints [ali2025environmental].
However, existing foundation models in Earth science focus primarily on gridded reanalysis data (weather/climate) or basin-scale streamflow prediction. To our knowledge, no prior work has applied foundation model principles to operational quality control for in situ sensor networks. This gap reflects unique challenges:
  • Extreme scale heterogeneity: Six orders of magnitude in discharge within a single dataset, requiring specialized normalization approaches
  • Irregular multi-site time series: Non-gridded data with site-specific characteristics (rating curves, instrument types, operational protocols)
  • Real-time deployment constraints: Inference must operate without access to ground truth, using only observable quantities
  • Deploy-safe requirements: The model must not corrupt valid data—a failure mode unacceptable in operational settings
  • Human oversight mandates: Agencies require interpretable outputs, uncertainty quantification, and audit trails for regulatory compliance
2.4 Positioning of HydroGEM
HydroGEM integrates advances from multiple domains while addressing the distinct challenges of operational hydrological quality control. From foundation models [nguyen2023climax, pathak2022fourcastnet], we adopt self-supervised pretraining on massive unlabeled datasets. From large-sample hydrology [kratzert2019towards, gauch2021proper], we use multi-site learning for cross-basin transfer. From recent deep learning work [dooley2023forecastpfn, kim2024enhancing], we use synthetic anomaly injection for task-specific fine-tuning. From operational QC systems [aquarius2024, schmidt2023saqc], we incorporate human-in-the-loop workflows and audit requirements.
The result is a system that: (1) learns from 3,505 diverse sites without site-specific tuning, (2) handles six orders of magnitude through hierarchical normalization, (3) detects diverse anomaly types via synthetic training, (4) shows rigorous zero-shot generalization through site-level evaluation, (5) provides suggested reconstructions for hydrologist review, and (6) operates deploy-safe with human oversight. While individual components draw inspiration from prior work, their integration in a validated operational system for continental-scale hydrological quality control represents a new contribution bridging machine learning and environmental monitoring.
@@ 2.4 Positioning of HydroGEM
Table 1 contrasts HydroGEM with related approaches.
Table 1: Positioning HydroGEM relative to related work
@@
Rule-based QC [aquarius2024, schmidt2023saqc] Any N/A N/A No Manual Yes
Single-site ML [he2019temporal, lee2024lstm] 1–10 Varies No Sometimes Varies No
Multi-basin hydro [kratzert2019towards] 100–500 No Yes No No N/A
Weather FM [nguyen2023climax, pathak2022fourcastnet] Gridded Yes Yes No No N/A
HydroGEM 5,421 Yes Yes Yes Suggested Yes

### v2

# HydroGEM: A Self-Supervised Zero Shot Hybrid TCN-Transformer Foundation Model for Continental Scale Streamflow Quality Control

@@ Abstract
Advances in sensor networks have enabled real-time, high-resolution stream discharge monitoring from in-situ gauges, yet persistent sensor malfunctions and data quality issues limit their utility. Manual quality control by expert hydrologists—while essential for maintaining data integrity—creates a bottleneck that delays scientific and operational use of these observations and cannot scale with networks generating millions of measurements annually. To address this gap, we introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a foundation model for continental-scale streamflow quality control designed to support human expertise rather than replace it. HydroGEM uses a two-stage training approach: self-supervised pretraining on 6.03 million clean sequences from 3,724 USGS stations learns general hydrological representations, followed by fine-tuning with synthetic anomalies for detection and reconstruction. A hybrid Temporal Convolutional Network–Transformer architecture (14.2M parameters) captures both local temporal patterns and long-range dependencies, while hierarchical normalization handles learning across six orders of magnitude in discharge. On a held-out test set of 799 stations with 18 expert-validated anomaly types, HydroGEM achieves F1 = 0.792 for detection and 68.7% reconstruction-error reduction, outperforming the strongest baseline by 36.3%. Zero-shot transfer to 100 Environment and Climate Change Canada stations—unseen during training—yields F1 = 0.586, demonstrating cross-national generalization. The model maintains consistent detection across correction magnitudes (1–100%) and aligns with operational seasonal patterns, with peak flagging rates during winter ice-affected periods matching hydrologists’ correction behavior. Architectural separation between simplified training anomalies and complex physical-space test anomalies confirms that performance reflects learned hydrometric principles rather than pattern memorization.
@@ 1 Introduction
Real-time hydrological monitoring networks provide essential observations for water resources management, flood forecasting, ecosystem protection, and climate change adaptation [28, 29, 50]. The United States Geological Survey operates more than 10,000 stream gauging stations that collectively produce millions of paired discharge and stage measurements each month, forming the backbone of decision-making across federal, state, and local agencies [27, 50]. High-frequency sensor data are only as useful as their quality, and without timely, accurate quality control, the growing volume of observations becomes a bottleneck rather than a resource [37].
Advances in sensor technology, telemetry, and network infrastructure have dramatically expanded the spatial and temporal density of hydrological monitoring. Modern gauging stations transmit stage and discharge at sub-hourly intervals, enabling near-real-time flood warnings, reservoir operations, and ecological flow management. Yet despite these technological improvements, raw sensor data remain inherently imperfect. Even in well-maintained monitoring networks, physical sensors inevitably experience drift, fouling, ice effects, rating-curve shifts, clock errors, and transmission failures, producing intermittent gaps or anomalies in the data [51, 12]. These issues are an inherent part of real-world sensing, where the central challenge is ensuring fast, reliable quality control of the data they generate.
Traditional quality control in hydrological monitoring combines manual expert review with rule-based automated checks [4, 47, 24]. Commercial systems such as AQUARIUS implement configurable range checks, rate-of-change thresholds, and statistical comparison filters, providing standardized workflows and audit trails [4]. Open-source frameworks like SaQC advance reproducibility through sequential test configurations and traceable quality flags [47]. However, rule-based approaches face fundamental limitations: fixed thresholds require site-specific calibration and capture only obvious outliers. Context-dependent anomalies—such as ice effects that preserve plausible values but bias discharge [35], gradual sensor drift within historical ranges [53], or rating curve shifts after channel-altering floods [1]—often evade detection. Multivariate relationships between discharge and stage, critical for identifying coupled sensor failures or unit conversion errors, are difficult to encode in rule-based systems. As monitoring networks expand and temporal resolution increases, manual calibration becomes prohibitively expensive.
Current operational practice therefore relies heavily on domain experts who visually inspect time series, interpret stage–discharge relationships, and apply site-specific judgment [37, 15]. While this expert-driven workflow produces high-quality records, its manual nature cannot keep pace with expanding monitoring networks and rising data frequency. The resulting bottleneck stems not from insufficient expertise but from the growing mismatch between data volumes and the capacity for timely human review.
Recent advances in machine learning offer a promising path toward more efficient and scalable hydrologic data quality control. Modern models can process continuous data streams, capture subtle temporal dependencies, and operate consistently across large station networks—capabilities well aligned with the needs of contemporary monitoring systems. In principle, such methods could provide rapid, reliable quality screening across thousands of sites.
Classical anomaly detection methods have been applied to environmental monitoring with mixed success. Statistical approaches (z-score tests, ARIMA residuals), distance-based methods (k-nearest neighbors, isolation forests), and one-class classifiers (one-class SVM, autoencoders) provide baseline capabilities [5, 49]. Recent applications include wavelet-based multi-resolution analysis for river stage [40] and comparative evaluations of isolation forests versus one-class SVM for groundwater monitoring, where OCSVM achieved 88% precision on synthetic data [49].
Deep learning approaches show stronger performance by capturing complex temporal dependencies. LSTM networks and Transformer attention mechanisms excel at modeling long-range patterns in time series [23, 21]. Temporal Convolutional Networks enable efficient parallel processing with exponentially large receptive fields [19]. Autoencoders—including variational and LSTM-augmented variants—detect anomalies via reconstruction errors [42, 31]. Generative Adversarial Networks improve robustness through learned representations of normal data distributions [8]. Synthetic anomaly generation has emerged as a key technique for training when domain-specific labeled data are unavailable [26].
@@
Translating this potential into practice, however, is challenging because continental-scale deployment faces two fundamental obstacles. First, hydrological systems are extremely heterogeneous: discharge spans six orders of magnitude across sites, and flow generation mechanisms vary with climate, physiography, and human modification [27, 29, 32]. Models typically train and evaluate on individual sites or small regional datasets (often fewer than 50 sites), failing to capture the extreme heterogeneity spanning snowmelt-dominated montane systems to ephemeral desert washes to regulated lowland rivers. Transfer learning across such diverse regimes remains largely unexplored in hydrological anomaly detection. Second, labeled anomaly datasets are scarce. Agency quality flags are not designed as ground truth labels that cleanly separate sensor faults from physical phenomena, and comprehensive labeled datasets across thousands of sites do not exist [12, 5].
Foundation models—large networks pretrained on massive datasets with self-supervised objectives, then transferred to diverse downstream tasks—provide a path forward [10, 13]. In weather and climate science, ClimaX demonstrated that transformers pretrained on heterogeneous climate reanalysis data with masked token prediction effectively transfer to forecasting, downscaling, and multi-variable prediction across spatiotemporal scales [38]. Weather foundation models including FourCastNet [43], GraphCast [30], and Pangu-Weather [7] now achieve medium-range forecast skill competitive with operational numerical weather prediction at orders of magnitude lower computational cost. Industrial-scale efforts such as IBM/NASA’s Prithvi WxC (2.3B parameters) [48] and Oak Ridge’s ORBIT (>>100B parameters) [58] are extending this paradigm to multi-variable Earth system prediction.
In hydrology, Kratzert and colleagues pioneered multi-basin learning by training LSTM networks across hundreds of catchments using the CAMELS dataset [27], demonstrating that models learn generalizable runoff generation processes that transfer to ungauged basins. Subsequent work confirmed that diverse training across basins improves predictions despite regional hydrological variation [14]. Remote sensing foundation models pretrained on satellite imagery have shown strong transfer to environmental monitoring tasks [2]. Recent perspectives advocate for foundation models as assistive tools in hydrometeorology when developed with appropriate domain constraints [2].
Yet foundation model principles have not been systematically applied to quality control for in situ sensor networks. This gap reflects unique challenges: extreme scale heterogeneity (six orders of magnitude in discharge within a single dataset, requiring specialized normalization); irregular multi-site time series (non-gridded data with site-specific rating curves, instrument types, and operational protocols); real-time deployment constraints (inference must operate without access to ground truth); deploy-safe requirements (the model must not corrupt valid data—a failure mode unacceptable in operational settings); and human oversight mandates (agencies require interpretable outputs, uncertainty quantification, and audit trails for regulatory compliance).
Our approach addresses these limitations by integrating advances from multiple domains. From foundation models [38, 43], we adopt self-supervised pretraining on massive unlabeled datasets. From large-sample hydrology [27, 14], we use multi-site learning for cross-basin transfer. From recent deep learning work [13, 26], we use synthetic anomaly injection for task-specific fine-tuning. From operational QC systems [4, 47], we incorporate human-in-the-loop workflows and audit requirements. The result is a system that learns multivariate patterns from continental-scale data, enabling context-aware detection without manual threshold tuning while preserving the human-in-the-loop workflows that agencies require.
@@
To address these challenges, we introduce HydroGEM (Hydrological Generalizable Encoder for Monitoring), a self-supervised foundation model for continental-scale streamflow quality control. HydroGEM qualifies as a foundation model because it: (1) trains on massive diverse data (3,724 sites spanning six orders of magnitude in discharge), (2) uses self-supervised pretraining to learn general hydrological representations, (3) shows zero-shot transfer to unseen sites and countries, and (4) uses a modular architecture that allows task-specific adaptation.
@@
HydroGEM uses a two-stage training approach. Stage 1 pretrains a hybrid TCN-Transformer backbone on 6.03 million clean sequences from 3,724 USGS sites using masked reconstruction. A hierarchical normalization scheme—combining log transforms, site-specific standardization, and explicit scale embeddings—allows learning across extreme magnitude ranges while preserving physically meaningful scale structure [17]. Stage 2 fine-tunes the pretrained backbone with a detection head using on-the-fly synthetic anomaly injection rather than curated labels [13, 26]. The training injector applies simplified corruptions (drift, spikes, flatlines, dropouts, clock shifts) in normalized space to encourage learning of fundamental hydrometric consistency principles rather than memorization of specific anomaly signatures. Clean data preservation exceeds 97%, ensuring the model does not corrupt valid observations.
@@
Evaluation emphasizes generalization across three dimensions. We construct a synthetic test set of 799 USGS sites (completely unseen during training) with expert-validated anomalies that are deliberately more complex than training corruptions, creating a four-axis separation (geographic, mathematical, temporal, parameter) that prevents pattern memorization. On this held-out set, HydroGEM achieves 0.70 F1 for anomaly detection and reduces reconstruction error by 70% relative to injected corruptions—though suggested corrections require expert review [24] before operational use. For real-world validation, we evaluate zero-shot transfer to 100 Environment and Climate Change Canada (ECCC) sites using weak labels derived from operational corrections. Reconstructions show promising alignment with agency corrections in several cases, though direct comparison is limited by differences in correction methodology between USGS and ECCC workflows. This cross-national evaluation tests whether the learned representations transfer beyond the training data source [27, 11].
@@
HydroGEM is designed for human-in-the-loop workflows: a three-tier flagging system passes high-confidence clean data with minimal review, prioritizes uncertain observations for expert inspection, and provides suggested reconstructions with uncertainty estimates. All outputs include provenance information suitable for audit and feedback, supporting hydrologist expertise [24, 37, 15].
@@
In summary, this work advances hydrological quality control through five contributions:
@@
    A continental-scale foundation model trained on 3,724 USGS sites with 6.03 million sequences—an order of magnitude larger than prior multi-site hydrological studies.
@@
Table 1 contrasts HydroGEM with related approaches across key dimensions.
Table 1: Positioning HydroGEM relative to related approaches
@@
Rule-based QC [4, 47] Any N/A N/A No Manual Yes
Single-site ML [21, 31] 1–10 Varies No Sometimes Varies No
Multi-basin hydro [27] 100–500 No Yes No No N/A
Weather FM [38, 43] Gridded Yes Yes No No N/A
HydroGEM 3,724 Yes Yes Yes Suggested Yes
Table LABEL:tab:terminology summarizes key technical terms used in this paper.*** Ijaz, can we refer to the terminology table as a LaTeX table here?
Terminology Clean data: USGS approved records with quality code ‘A’ (verified and approved by USGS hydrologists as accurate representations of observed conditions; used for Stage 1 training) Masked data: Clean data with random timesteps hidden for self-supervised reconstruction (Stage 1) Corrupted data: Clean data with synthetic anomalies injected (Stage 2 fine-tuning) Anomalous data: Real operational data requiring correction (deployment target) Suggested reconstruction: Model-proposed correction requiring hydrologist approval Stage/Gage height: Water surface elevation measured at the gauging station, typically in feet Discharge: Volume of water flowing past a point per unit time, typically in cubic feet per second (ft3/s) Rating curve: Empirical relationship between stage and discharge at a specific site